OpenStudy (anonymous):

Solve: `x - 17 = 23` A. x = 6 B. x = 40 C. x = -40 D. x = -6

OpenStudy (anonymous):

\[x-17=23\] \[x-17+17=23+17\] \[x=40\] Therefore C is the answer